The amount of money that a salesperson makes varies directly as the total amount of sales made. If the salesperson receives $7,500 for selling a house for $120,000, how much will he or she make if he or she sells a house for $150,000?
Answer by scott8148(6628)

let x=comission

7500/120000=x/150000 __ 7500/12=x/15 __ 37500/4=x __ 9375=x